OpenStudy (anonymous):

Express the ratio in simplest form. 24 : 18 A. 4 : 3 B. 8 : 6 C. 6 : 1 D. 12 : 9

OpenStudy (anonymous):

Divide each of the number by their hcf

OpenStudy (anonymous):

hcf of of 24 and 18=6 therefore 24:18 in its simplest form=24/6:18/6=??? @theoryofmars

OpenStudy (anonymous):

@Curious9876 wouldnt it be 4/3 is the simplified form be because 24:18 is basically a fraction

OpenStudy (anonymous):

yes 4:3=4/3

OpenStudy (anonymous):

@ theoryofmars Think about the ratio 18:24 24 and 18 both have the factor 6. 24 divided by 6 = 4 18 divided by 6 = 3 So 24:18 is the same as 4:3

OpenStudy (anonymous):

and @Curious9876 okaii thats what i thought

OpenStudy (anonymous):

Yes,I meant the same.The GCD(HCF) of the numbers=6 so divide the numbers by 6.u l get 4:3 @$boricua

OpenStudy (anonymous):

i hope we both helped you

OpenStudy (anonymous):

thank you i actually had this answer i was just double checking ,sorry for not telling you guys sooner i had to help my brother with studyisland stuff !!

OpenStudy (anonymous):

is okay don't worry about it :)

OpenStudy (anonymous):

ook @theoryofmars i m happy that i cud help u

OpenStudy (anonymous):

:)

OpenStudy (haseeb96):

Sorry mujhe nahi malm tha k apko WO acha nahi lage ga@shereenkhan